OpenStudy (anonymous):

Consider the functions f(x) = 4x2 and g(x) = 5(x + 7). Find f[g(−5)]

OpenStudy (anonymous):

that was supposed to be" Consider the functions f(x) = 4x^2 and g(x) = 5(x + 7). Find f[g(−5)]"

OpenStudy (anonymous):

Consider the functions f(x) = 4x^2 and g(x) = 5(x + 7). Find f[g(−5)]"